# Random Spiking Neural Networks are Stable and Spectrally Simple

### v1

@@ 1 Introduction
Artificial Neural Networks (ANNs) have become central to modern machine learning, but their rapid growth in scale demands increasingly unsustainable computational and energy resources (Thompson et al., 2021). This challenge is especially acute for low-power devices, where efficiency is critical. Spiking Neural Networks (SNNs), inspired by biological neurons and operating through event-driven spikes, offer a promising energy-efficient alternative (Roy et al., 2019; Davies et al., 2018). Their sparse communication and compatibility with neuromorphic hardware position them as strong candidates for sustainable machine learning (Mehonic et al., 2024; Fono et al., 2025).
@@
Despite this potential, the theoretical understanding of SNNs remains limited compared to that of classical ANNs. While substantial progress has been made on designing training algorithms (Eshraghian et al., 2023) and hardware implementations (Davies et al., 2018; Indiveri & Liu, 2015), core theoretical properties—such as stability, robustness, and generalization—are still largely underexplored. This gap limits our ability to rigorously assess the strengths and limitations of SNNs in practice.
@@
Among these properties, stability is crucial for designing networks resilient to input perturbations and adversarial attacks, yet no single definition exists. It can refer to algorithmic stability of learning algorithms (Elisseeff et al., 2005), dynamical systems stability (naturally aligned with the temporal dynamics of spiking neurons) (Ding et al., 2024), or—as we consider here—to sensitivity to input changes. Intuitively, a stable network should be resilient to small perturbations in inputs or parameters, which is critical for both reliable inference and efficient learning. Many SNN models, including the discrete-time Leaky Integrate-and-Fire (LIF) neuron, can be viewed as iterative compositions of Boolean functions, since neurons emit spikes only when their membrane potential crosses a threshold (a binary event), motivating the study of stability via Boolean function analysis (O’Donnell, 2014). While prior work has examined SNN stability from dynamical systems (Ding et al., 2024) or neuroscience perspectives (Calaim et al., 2022), to our knowledge this is the first study to apply Boolean analysis to characterize and investigate stability in SNNs.
@@
The Boolean perspective on SNNs suggests a natural link to the notion of simplicity in neural networks: if a network is stable, small input perturbations rarely change its output, hinting at a bias toward “simple” input-output mappings. In the SNN setting, this can be formalized via the predominance of low-frequency components in the Fourier–Walsh expansion of a classifier. Motivated by this, we introduce spectral simplicity, which quantifies the concentration of the Fourier spectrum on low frequencies. This notion connects naturally with other notions of simplicity studied in the context of simplicity bias, proposed as a lens to understand generalization in deep networks. For example, prior work has shown that random deep ANNs tend to implement “simple” functions (De Palma et al., 2019), formalized as a large average Hamming distance to the nearest input with a different predicted class (Valle-Pérez et al., 2019). Spectral simplicity represents a weaker notion, but one that arises intrinsically in spiking networks.
@@
To make these questions precise, we focus on the discrete-time LIF model. This model combines theoretical simplicity with widespread adoption in practice, serving as the foundation of software frameworks such as SNN-Torch and being implemented in digital neuromorphic platforms including Loihi 2 and SpiNNaker 2 (Orchard et al., 2021; Gonzalez et al., 2023). In this work, we restrict attention to networks at initialization. This choice reflects the still-developing theory of SNN training and our goal of isolating stability properties intrinsic to the model, without confounding effects from learning dynamics. Moreover, random networks have been shown to serve as useful priors in PAC-Bayes generalization bounds (Valle-Pérez et al., 2019). Understanding the effect of training dynamics is an exciting question; here, we explore it experimentally and leave a theoretical treatment for future work.
@@
  • We derive quantitative bounds on the stability of discrete-time LIF SNN classifiers, showing that they are stable on average with respect to random parameter initialization. In particular, when input sequences lie in the binary cube $\{-1,1\}^{n}$, the classifier output remains unchanged under perturbations of up to $\mathcal{O}(\sqrt{n})$ coordinates, with high probability for $n$ large enough. The bounds depend explicitly on the model’s hyperparameters and reveal that LIF SNNs exhibit stability properties comparable to other (time-independent) Boolean networks (Jonasson et al., 2023).
@@ 1.1 Related works
As previously noted, stability in SNNs can be approached from multiple perspectives. Each neuron follows a (typically non-linear) dynamical system, raising the natural question of how input perturbations affect its output. In the ANN and Neural ODE literature, Lyapunov-based analyses (Jimenez-Rodriguez et al., 2022; Kang et al., 2021; Rahnama et al., 2019) establish robustness via bounds on how perturbations propagate through the network. The discrete-time LIF model’s robustness to input perturbations has been analyzed from a dynamical systems perspective by Ding et al. (2024), who derived bounds on output spike sequence variations under perturbed inputs. Our work differs in three key ways: (i) we study stability at the classifier level, where predictions may remain unchanged even if spike trains differ; (ii) we consider the reset-by-subtraction mechanism, which introduces additional complexity compared to the reset-to-zero simplification in (Ding et al., 2024); and (iii) our analysis extends beyond single neurons to multi-neuron networks. Methodologically, our approach is different building on Boolean function analysis (O’Donnell, 2014), aligning more closely with Jonasson et al. (2023), but introduces new challenges stemming from reset dynamics and temporal evolution, which create nontrivial probabilistic dependencies.
@@
Simplicity bias has been proposed as a mechanism underlying the generalization of deep networks, both in trained models with SGD (Arpit et al., 2017; Nakkiran et al., 2019; Valle-Pérez et al., 2019) and in random ANNs (De Palma et al., 2019). The central idea is that learning favors simple functions, though the precise definition of simplicity varies, and pitfalls have been noted in (Shah et al., 2020). We introduce spectral simplicity, based on the Fourier–Walsh decomposition of Boolean functions, as a complementary notion. While our focus on random networks parallels (De Palma et al., 2019), the techniques differ substantially: their analysis relies on Gaussian process arguments, whose extension to SNNs is unclear, whereas we draw on Boolean function analysis, which adapts naturally to our setting. While a weaker measure of simplicity, spectral simplicity arises organically in spiking networks.
@@ 2 The Discrete-Time LIF Model
Each neuron processes an input sequence $(x_{t})_{t\in[T]}$ through its membrane potentials $(u_{t})_{t\in[T]}$, which evolve according to an autoregressive decay dynamics over the time horizon $T$ (referred to as the latency of the model). The leak parameter $\beta\in[0,1]$ controls this decay, specifying the fraction of potential retained per time step. Whenever the membrane potential exceeds the activation threshold $\theta>0$ at some time $t^{\prime}$, the neuron emits a spike, recorded as $s_{t^{\prime}}=1$ in the spike train $(s_{t})_{t\in[T]}$, and the potential is reduced by $\theta$ (reset by subtraction). Weights are initialized as $w\sim\mathcal{N}(0,I_{n}/n)$, ensuring $w^{\top}x=O(1)$ with high probability for $x\in\{-1,1\}^{n}$, thereby avoiding degenerate regimes of vanishing or overly frequent firing. For further background on the LIF model, see (Gerstner & Kistler, 2002). In short, we use the term leaky integrate-and-fire (LIF) neuron for the variant with a Heaviside step function and reset rule $v_{t}\mapsto v_{t}-\theta s_{t-1}$, and integrate-and-fire (IF) when the leakage is absent ($\beta=1$); networks composed of such units are referred to as LIF and IF SNNs, respectively.
@@
Given a $L$-layers sLIF SNN as in Definition 2, a widely used choice of classifier in SNNs (see e.g., (Diehl & Cook, 2015)) is based on spike counts at the output layer, namely
@@
Throughout this work, we assume $n_{0}=\dots=n_{L-1}=n$, with $n_{L}$ equal to the number of classes. For simplicity, we focus on a variant of the sLIF model equation 2 with $\beta=1$. Our analysis allows dynamic input sequences $(x_{t})_{t\in[T]}$, though some results only apply to static inputs, interpreted as repeated presentations of the same sample over time. Such constant input encoding is commonly used in practice for time-static datasets such as MNIST or CIFAR-10 (Rathi & Roy, 2023; Rueckauer et al., 2017). Finally, the model parameters are initialized as
@@ 3 Noise sensitivity of boolean functions
Within this framework, sLIF neurons can be viewed as compositions of Boolean functions, allowing stability analysis via Boolean function theory. We recall the basic definitions here and refer the reader to (O’Donnell, 2014) for a comprehensive treatment. Additionally, we introduce spectral simplicity, which is central to the statement of our main result.
@@
Every $f:\{-1,1\}^{n}\to\mathbb{R}$ admits a unique Fourier–Walsh expansion (O’Donnell, 2014, Thm. 1.1):
@@
Low-degree terms ($|S|$ small) correspond to low frequencies, and high-degree terms to high frequencies. A constant function, for instance, has spectrum supported on $\emptyset$. The notion of spectrum concentration formalizes when a function is “simple”: most of its Fourier weight lies on low-degree terms, i.e., subsets of size $o(n)$. Formally (O’Donnell, 2014, Def. 3.1), $f$ is $\epsilon$-concentrated up to degree $k$ if
@@
A key connection between noise stability and spectral concentration is given by (O’Donnell, 2014, Prop. 3.3), which states that for a Boolean function $f$, if we set $\epsilon=3\,\operatorname{\bf{NS}}_{\nu}(f)$, then the spectrum of $f$ is $\epsilon$-concentrated up to degree $1/\nu$. This result extends naturally to parametric families of functions by linearity, replacing $\operatorname{\bf{NS}}_{\nu}(f)$ with $\operatorname{\bf{ENS}}_{\nu}(f)$ and spectral concentration with its expected counterpart (see Lemma 6).
@@
Alternative notions of simplicity have been proposed in the simplicity bias literature (Valle-Pérez et al., 2019; De Palma et al., 2019). Closest to our setting is the definition of De Palma et al. (2019), who study, for a classifier $f:\{-1,1\}^{n}\to\{-1,1\}$ and $x\in\{-1,1\}^{n}$, the quantity
@@ 4 Noise stability of SNN classifiers
We proceed by induction on $l\in[L]$. To illustrate, consider $t=1$. Following (Jonasson et al., 2023), we analyze the Markov chain
@@
To illustrate, take $\nu^{\prime}=\tfrac{1}{\sqrt{n\log n}}$. In this case, an $L$-layer binary SNN classifier is $\mathcal{O}\!\big(n^{1/{2^{2(L+1)}}}\big)$-concentrated up to degree $\mathcal{O}(\sqrt{n\log n})$. Thus, only a vanishing fraction of degrees contribute meaningfully to the spectrum, making these classifiers spectrally simple. Interestingly, the bound on the maximal degree of concentration is independent of architectural parameters, while the concentration level deteriorates with larger $L$, $T$, and $\theta$ (the $\log^{3/2}n$ seems an artifact of the proof). The increase with $L$ is not surprising, since compositions of Boolean networks typically behave similarly, and analogous results are known for threshold ANNs (Jonasson et al., 2023). The $\theta$-dependence appears to be a proof artifact. Whether the $T$- and $L$-dependencies are intrinsic remains an open question, which we investigate experimentally in the next section.
@@ 5 Numerical experiments
We evaluate the noise sensitivity of trained sIF and IF SNNs in static data. In particular, we train three-layer sLIF and IF SNN on MNIST (i.e., $n=784$) using the ADAM optimizer with surrogate gradients (Eshraghian et al., 2023) until reaching $98\%$ training accuracy. After training, we estimate the output sensitivity by perturbing each test sample $100$ times, flipping each component with probability $\nu\in\left[\tfrac{1}{n},\tfrac{2}{n}\dots,\tfrac{1}{\sqrt{n}}\right]$ and measuring the fraction of perturbations that change the network’s output. For comparison with a random network, the noise sensitivity is estimated using the same procedure, averaged over $10$ independent random weight initializations. As expected, training significantly reduces the sensitivity of the model whenever the final test accuracy is sufficiently high. Figure 3(a) shows show the results for MNIST. Figure 3 shows the same experiment on CIFAR-10 ($n=3072$). Notice that training reduces the model’s sensitivity, but less strongly compared to MNIST. This aligns with the fact that both the training and test errors are larger for CIFAR-10 (which achieves $84.38\%$ training accuracy).
@@ 6 Conclusion
  • Investigating the average distance to the nearest input with a different label (see equation 4), as in (De Palma et al., 2019). Unlike their setting, a simple union bound fails here, since our $\mathcal{O}(1/\sqrt{n})$ bound is insufficient given the exponentially many inputs at distance $\mathcal{O}(\sqrt{n})$.
@@ Appendix B Technical Lemmas
We now give a straightforward generalization of (O’Donnell, 2014, Prop.3.3.) to the case of expected spectrum concentration (Definition 4).
@@
Following (O’Donnell, 2014), for a fixed $f_{w}$,
@@
where the $\mathbf{S}_{f_{w}}$ denotes the probability distribution over the subsets of $[n]$, which assigns probability $\hat{f}_{w}^{2}(S)$ to the subset $S$ (recall that the Fourier coefficients sum to $1$). Then, using (O’Donnell, 2014, Thm.2.49)

### v2

@@ 1 Introduction
Artificial Neural Networks (ANNs) have become central to modern machine learning, but their rapid growth in scale demands increasingly unsustainable computational and energy resources (DL_diminishing_returns_2021). This challenge is especially acute for low-power devices, where efficiency is critical. Spiking Neural Networks (SNNs), inspired by biological neurons and operating through event-driven spikes, offer a promising energy-efficient alternative (roy2019towards; davies2018loihi). Their sparse communication and compatibility with neuromorphic hardware position them as strong candidates for sustainable machine learning (Mehonic2024NeuromorphicRoadmap; fono2025sustainableaimathematicalfoundations).
@@
Despite this potential, the theoretical understanding of SNNs remains limited compared to that of classical ANNs. While substantial progress has been made on designing training algorithms (Lessons_from_DL_eshraghian_2023) and hardware implementations (davies2018loihi; indiveri2015memory), core theoretical properties—such as stability, robustness, and generalization—are still largely underexplored. This gap limits our ability to rigorously assess the strengths and limitations of SNNs in practice.
@@
Among these properties, stability is crucial for designing networks resilient to input perturbations and adversarial attacks, yet no single definition exists. It can refer to algorithmic stability of learning algorithms (Algo_stability_1), dynamical systems stability (naturally aligned with the temporal dynamics of spiking neurons) (Robust_SNN), or—as we consider here—to sensitivity to input changes. Intuitively, a stable network should be resilient to small perturbations in inputs or parameters, which is critical for both reliable inference and efficient learning. Many SNN models, including the discrete-time Leaky Integrate-and-Fire (LIF) neuron, can be viewed as iterative compositions of Boolean functions, since neurons emit spikes only when their membrane potential crosses a threshold (a binary event), motivating the study of stability via Boolean function analysis (Odonnell). While prior work has examined SNN stability from dynamical systems (Robust_SNN) or neuroscience perspectives (Robustness_SNN_neuro), to our knowledge this is the first study to apply Boolean analysis to characterize and investigate stability in SNNs.
@@
The Boolean perspective on SNNs suggests a natural link to the notion of simplicity in neural networks: if a network is stable, small input perturbations rarely change its output, hinting at a bias toward “simple” input-output mappings. In the SNN setting, this can be formalized via the predominance of low-frequency components in the Fourier–Walsh expansion of a classifier. Motivated by this, we introduce spectral simplicity, which quantifies the concentration of the Fourier spectrum on low frequencies. This notion connects naturally with other notions of simplicity studied in the context of simplicity bias, proposed as a lens to understand generalization in deep networks. For example, prior work has shown that random deep ANNs tend to implement “simple” functions (depalma2019random), formalized as a large average Hamming distance to the nearest input with a different predicted class (VallePerez2019). Spectral simplicity represents a weaker notion, but one that arises intrinsically in spiking networks.
@@
To make these questions precise, we focus on the discrete-time LIF model. This model combines theoretical simplicity with widespread adoption in practice, serving as the foundation of software frameworks such as SNN-Torch and being implemented in digital neuromorphic platforms including Loihi 2 and SpiNNaker 2 (Loihi2_2021; spinnaker_async_ML_2024). In this work, we restrict attention to networks at initialization. This choice reflects the still-developing theory of SNN training and our goal of isolating stability properties intrinsic to the model, without confounding effects from learning dynamics. Moreover, random networks have been shown to serve as useful priors in PAC-Bayes generalization bounds (VallePerez2019). Understanding the effect of training dynamics is an exciting question; here, we explore it experimentally and leave a theoretical treatment for future work.
@@
  • We derive quantitative bounds on the stability of discrete-time LIF SNN classifiers, showing that they are stable on average with respect to random parameter initialization. In particular, when input sequences lie in the binary cube $\{-1,1\}^{n}$, the classifier output remains unchanged under perturbations of up to $\mathcal{O}(\sqrt{n})$ coordinates, with high probability for $n$ large enough. The bounds depend explicitly on the model’s hyperparameters and reveal that LIF SNNs exhibit stability properties comparable to other (time-independent) Boolean networks (jonasson2023noise).
@@ 1.1 Related works
As previously noted, stability in SNNs can be approached from multiple perspectives. Each neuron follows a (typically non-linear) dynamical system, raising the natural question of how input perturbations affect its output. In the ANN and Neural ODE literature, Lyapunov-based analyses (jimenez2022lyanet; kang2021stable; rahnama2019robust) establish robustness via bounds on how perturbations propagate through the network. The discrete-time LIF model’s robustness to input perturbations has been analyzed from a dynamical systems perspective by Robust_SNN, who derived bounds on output spike sequence variations under perturbed inputs. Our work differs in three key ways: (i) we study stability at the classifier level, where predictions may remain unchanged even if spike trains differ; (ii) we consider the reset-by-subtraction mechanism, which introduces additional complexity compared to the reset-to-zero simplification in (Robust_SNN); and (iii) our analysis extends beyond single neurons to multi-neuron networks. Methodologically, our approach is different building on Boolean function analysis (Odonnell), aligning more closely with jonasson2023noise, but introduces new challenges stemming from reset dynamics and temporal evolution, which create nontrivial probabilistic dependencies.
@@
Simplicity bias has been proposed as a mechanism underlying the generalization of deep networks, both in trained models with SGD (arpit2017closer; nakkiran2019sgd; VallePerez2019) and in random ANNs (depalma2019random). The central idea is that learning favors simple functions, though the precise definition of simplicity varies, and pitfalls have been noted in (pitfalls_SB2020). We introduce spectral simplicity, based on the Fourier–Walsh decomposition of Boolean functions, as a complementary notion. While our focus on random networks parallels (depalma2019random), the techniques differ substantially: their analysis relies on Gaussian process arguments, whose extension to SNNs is unclear, whereas we draw on Boolean function analysis, which adapts naturally to our setting. While a weaker measure of simplicity, spectral simplicity arises organically in spiking networks.
@@ 2 The Discrete-Time LIF Model
Each neuron processes an input sequence $(x_{t})_{t\in[T]}$ through its membrane potentials $(u_{t})_{t\in[T]}$, which evolve according to an autoregressive decay dynamics over the time horizon $T$ (referred to as the latency of the model). The leak parameter $\beta\in[0,1]$ controls this decay, specifying the fraction of potential retained per time step. Whenever the membrane potential exceeds the activation threshold $\theta>0$ at some time $t^{\prime}$, the neuron emits a spike, recorded as $s_{t^{\prime}}=1$ in the spike train $(s_{t})_{t\in[T]}$, and the potential is reduced by $\theta$ (reset by subtraction). Weights are initialized as $w\sim\mathcal{N}(0,I_{n}/n)$, ensuring $w^{\top}x=O(1)$ with high probability for $x\in\{-1,1\}^{n}$, thereby avoiding degenerate regimes of vanishing or overly frequent firing. For further background on the LIF model, see (gerstner2002spiking). In short, we use the term leaky integrate-and-fire (LIF) neuron for the variant with a Heaviside step function and reset rule $v_{t}\mapsto v_{t}-\theta s_{t-1}$, and integrate-and-fire (IF) when the leakage is absent ($\beta=1$); networks composed of such units are referred to as LIF and IF SNNs, respectively.
@@
Given a $L$-layers sLIF SNN as in Definition 2, a widely used choice of classifier in SNNs (see e.g., (digit_recog_SNN)) is based on spike counts at the output layer, namely
@@
Throughout this work, we assume $n_{0}=\dots=n_{L-1}=n$, with $n_{L}$ equal to the number of classes. For simplicity, we focus on a variant of the sLIF model equation 2 with $\beta=1$. Our analysis allows dynamic input sequences $(x_{t})_{t\in[T]}$, though some results only apply to static inputs, interpreted as repeated presentations of the same sample over time. Such constant input encoding is commonly used in practice for time-static datasets such as MNIST or CIFAR-10 (DIET-SNN_direct_input_encoding_2023; direct_input_2_rueckauer_2017). Finally, the model parameters are initialized as
@@ 3 Noise sensitivity of boolean functions
Within this framework, sLIF neurons can be viewed as compositions of Boolean functions, allowing stability analysis via Boolean function theory. We recall the basic definitions here and refer the reader to (Odonnell) for a comprehensive treatment. Additionally, we introduce spectral simplicity, which is central to the statement of our main result.
@@
Every $f:\{-1,1\}^{n}\to\mathbb{R}$ admits a unique Fourier–Walsh expansion (Odonnell, Thm. 1.1):
@@
Low-degree terms ($|S|$ small) correspond to low frequencies, and high-degree terms to high frequencies. A constant function, for instance, has spectrum supported on $\emptyset$. The notion of spectrum concentration formalizes when a function is “simple”: most of its Fourier weight lies on low-degree terms, i.e., subsets of size $o(n)$. Formally (Odonnell, Def. 3.1), $f$ is $\epsilon$-concentrated up to degree $k$ if
@@
A key connection between noise stability and spectral concentration is given by (Odonnell, Prop. 3.3), which states that for a Boolean function $f$, if we set $\epsilon=3\,\operatorname{\bf{NS}}_{\nu}(f)$, then the spectrum of $f$ is $\epsilon$-concentrated up to degree $1/\nu$. This result extends naturally to parametric families of functions by linearity, replacing $\operatorname{\bf{NS}}_{\nu}(f)$ with $\operatorname{\bf{ENS}}_{\nu}(f)$ and spectral concentration with its expected counterpart (see Lemma 6).
@@
Alternative notions of simplicity have been proposed in the simplicity bias literature (VallePerez2019; depalma2019random). Closest to our setting is the definition of depalma2019random, who study, for a classifier $f:\{-1,1\}^{n}\to\{-1,1\}$ and $x\in\{-1,1\}^{n}$, the quantity
@@ 4 Noise stability of SNN classifiers
We proceed by induction on $l\in[L]$. To illustrate, consider $t=1$. Following (jonasson2023noise), we analyze the Markov chain
@@
To illustrate, take $\nu^{\prime}=\tfrac{1}{\sqrt{n\log n}}$. In this case, an $L$-layer binary SNN classifier is $\mathcal{O}\!\big(n^{1/{2^{2(L+1)}}}\big)$-concentrated up to degree $\mathcal{O}(\sqrt{n\log n})$. Thus, only a vanishing fraction of degrees contribute meaningfully to the spectrum, making these classifiers spectrally simple. Interestingly, the bound on the maximal degree of concentration is independent of architectural parameters, while the concentration level deteriorates with larger $L$, $T$, and $\theta$ (the $\log^{3/2}n$ seems an artifact of the proof). The increase with $L$ is not surprising, since compositions of Boolean networks typically behave similarly, and analogous results are known for threshold ANNs (jonasson2023noise). The $\theta$-dependence appears to be a proof artifact. Whether the $T$- and $L$-dependencies are intrinsic remains an open question, which we investigate experimentally in the next section.
@@ 5 Numerical experiments
We evaluate the noise sensitivity of trained sIF and IF SNNs in static data. In particular, we train three-layer sLIF and IF SNN on MNIST (i.e., $n=784$) using the ADAM optimizer with surrogate gradients (Lessons_from_DL_eshraghian_2023) until reaching $98\%$ training accuracy. After training, we estimate the output sensitivity by perturbing each test sample $100$ times, flipping each component with probability $\nu\in\left[\tfrac{1}{n},\tfrac{2}{n}\dots,\tfrac{1}{\sqrt{n}}\right]$ and measuring the fraction of perturbations that change the network’s output. For comparison with a random network, the noise sensitivity is estimated using the same procedure, averaged over $10$ independent random weight initializations. As expected, training significantly reduces the sensitivity of the model whenever the final test accuracy is sufficiently high. Figure 3(a) shows show the results for MNIST. Figure 3 shows the same experiment on CIFAR-10 ($n=3072$). Notice that training reduces the model’s sensitivity, but less strongly compared to MNIST. This aligns with the fact that both the training and test errors are larger for CIFAR-10 (which achieves $84.38\%$ training accuracy).
@@ 6 Conclusion
  • Investigating the average distance to the nearest input with a different label (see equation 4), as in (depalma2019random). Unlike their setting, a simple union bound fails here, since our $\mathcal{O}(1/\sqrt{n})$ bound is insufficient given the exponentially many inputs at distance $\mathcal{O}(\sqrt{n})$.
@@ Appendix B Technical Lemmas
We now give a straightforward generalization of (Odonnell, Prop.3.3.) to the case of expected spectrum concentration (Definition 4).
@@
Following (Odonnell), for a fixed $f_{w}$,
@@
where the $\mathbf{S}_{f_{w}}$ denotes the probability distribution over the subsets of $[n]$, which assigns probability $\hat{f}_{w}^{2}(S)$ to the subset $S$ (recall that the Fourier coefficients sum to $1$). Then, using (Odonnell, Thm.2.49)